Home  Contact  Equipment  Field Notes  Proc Notes  Photos  Data  Logs Today's Date:   
   
   
  Biomet public Home
Active Sites
 BA 2023-09-28 - DS
 BC 2023-09-28 - DS
 EE 2023-09-28 - DS
 GT 2023-09-28 - DS
 HS 2023-09-28 - DS
 MB 2023-09-28 - DS
 SW 2022-11-18 - DS
 TZ 2023-09-19 - JV
 VR 2023-09-19 - JV
 WP 2023-09-28 - DS

Recently Added Notes
 GT 2023-09-28 - DS
 HS 2023-09-28 - DS
 WP 2023-09-28 - DS
 MB 2023-09-28 - DS
 EE 2023-09-28 - DS

   
   
   
   
   
 
 
<--2023-05-17 12:55:00 | 2023-06-14 10:30:00-->
Other sites visited today: Gilbert Tract | Mayberry | West Pond | East End | Bouldin Alfalfa | Hill Slough

Bouldin Corn: 2023-05-31 15:55:00 (DOY 151)
Author: Daphne Szutu
Others: Carlos, Jerry

Summary: Regular data collection/cleaning, baby corn!, installed zcams, Charlotte working on terrestrial lidar survey

2023-05-31
Bouldin Corn

Carlos, Jerry Sun, and I arrived at 15:55 PDT. It's been a cool spring so it was a relatively pleasant day, mild and clear but pretty windy. Jerry is an undergrad student exploring ideas for his senior thesis and today was his first time in the field.

Baby corn is growing!

Charlotte has been here all day doing her terrestrial lidar survey and has a few more hours to go. She joked her thesis will be called "When the Spheres Align" because of all her effort to align the white spherical sensors for the lidar.

I set up the 2 undercanopy zcams. The power cable seemed long in the lab but is a little short. There's a ~2m gap between the tower and the corn, so the first camera is right under the first row. Not ideal because it will see the edge of the field. I'll bring a longer power cable to swap out in two weeks. Even though I set both of the clocks in the lab, one of the cameras needed its clock set. Next time, I need to check the RTCs are attached securely, and add medium desiccant packs to the cameras.

I downloaded met, USB GHG, PA, and IRT data. No camera data again for a simple reason--the camera cable was not plugged into the datalogger. Fixed and confirmed that new photos were getting written to the CF card. Carlos cleaned flux and rad sensors, topped off the wash reservoir, and took veg height measurements. I changed the eddy clock +45sec to match the laptop time and restarted logging.

The 7500 read: 358ppm CO2, 959mmol/m3 H2O, 25.0C, 100.9kPa, 92SS--101 after cleaning
The 7700 read: 1.95ppm CH4, 33RSSI--81 after cleaning

Charlotte and Carlos chatted while I took pictures of the Duck Club breaker box for Joe. He is planning to add more power for Charlotte.

We left at 16:40 PDT.

Veg Height (cm): Tape measure from ground to top of plant
9, 11, 8, 11, 7, 8, 11, 7, 14, 13, 8, 10, 9, 11, 12, avg = 9.9, stdev = 2.15

0 photos found

Graphs display: days
6 sets found

17 BC_met graphs found

5 BC_flux graphs found

Can't check TATWBABCSB_cam
1 TATWBABCSB_cam graphs found

Can't check MBSWEEWPBCBA
1 MBSWEEWPBCBA graphs found

2 BC_pa graphs found

1 BC_arable graphs found

 
 

Login
Select Notes:

Search Title for:

Search Descritption for:

This material is based upon work supported by the National Science Foundation and US Department of Energy. Any opinions, findings, conclusions, or recommendations expressed in the material are those of the author(s) and do not necessarily reflect the views of the supporters.